OpenStudy (anonymous):

Choose the correct simplification of the expression 9/h^-3

OpenStudy (jdoe0001):

\(\bf \cfrac{1}{a^{-\color{red} n}}\implies \cfrac{1}{\frac{1}{a^{\color{red} n}}}\implies a^{{\color{red} n}} \\ \quad \\ \quad \\ \cfrac{9}{h^{-{\color{red}{ 3}}}}\implies \cfrac{9}{1}\cdot \cfrac{1}{h^{-{\color{red}{ 3}}}}\implies 9\cdot \square ?\)

OpenStudy (anonymous):

is the answer 9h^3 ?

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

thanks :)

OpenStudy (jdoe0001):

yw